# On the Origin of the High Star-Formation Efficiency in Massive Galaxies at Cosmic Dawn

### v1

@@ Appendix E The density PDF on different scales
We also convolve the density distribution with the portion of the PDF that is gravitationally unstable i.e. $\alpha_{\rm vir}<1$, effectively computing the density distribution of collapsing clouds. The tail of this distribution crosses the opacity limit $\rho\sim 10^{-13}~{}{\rm g/cm^{3}}$ (Masunaga & Inutsuka, 1999), where our model of a cold, isothermal gas begins to break down.
@@ Appendix G Column density PDF
The volume density PDF (Eq. 10) described by the MFF model can be related to the column density PDF, which can be directly measured from observations. Numerical simulations show that the column density PDF of supersonic turbulent gas is well-described by a log-normal distribution, similar to the volume density PDF (Federrath et al., 2010)
$p(\eta)=\frac{1}{\sqrt{2\pi\sigma_{\eta}^{2}}}\exp\left[-\frac{(\eta-\overline{\eta})^{2}}{2\sigma_{\eta}^{2}}\right]$ (68)
where $\eta=\ln(\Sigma/\overline{\Sigma})$ is the logarithmic surface density and $\sigma_{\eta}$ is the standard deviation of the logarithmic surface density in analogy to $s$ and $\sigma_{s}$ for the volume density PDF. We use $\eta$ to denote the logarithmic surface density for notational consistency with other work and rely on context to disambiguate the logarithmic surface density and outflow efficiency.
@@
Using that relationship, Seon (2012) show that the relationship between the two dispersions is well-approximated by a constant of proportionality $\sigma_{\eta}^{2}=A\sigma_{s}^{2}$, where $A$ depends only on the forcing parameter. For $b_{\rm turb}=1/3,0.5,1$, they find best fit values $A=0.2,0.24,0.38$ respectively.
We fit these data with a linear function using the least-squares method, which combined with Equation 12, yields
@@
Given the column density PDF $p(\Sigma)=p(\eta)/\Sigma$ and assuming a constant opacity $\kappa$, the optical depth PDF is simply $p(\tau)=\kappa p(\Sigma)$. The effective optical depth for a turbulent region can be estimated using a Rosseland-like averaging, which is equivalent to a PDF-weighted average over photon mean free path:
@@
When applied to a cell in our simulations of size $\Delta x$, the mean surface density is $\overline{\Sigma}=\overline{\rho}\Delta x$.

### v2

@@ Appendix E The density PDF on different scales
We also convolve the density distribution with the portion of the PDF that is gravitationally unstable i.e. $\alpha_{\rm vir}<1$, effectively computing the density distribution of collapsing clouds. The tail of this distribution crosses the opacity limit $\rho\sim 10^{-13}~{}{\rm g/cm^{3}}$ (Masunaga & Inutsuka, 1999), where the assumption of a cold, isothermal gas begins to break down.
@@ Appendix G Column density PDF
The volume density PDF (Eq. 10) described by the MFF model is related to the column density PDF, which is a direct observable. Numerical simulations show that the column density PDF of supersonic turbulent gas is well-described by a log-normal distribution, similar to the volume density PDF (Federrath et al., 2010)
[…]
where $\eta=\ln(\Sigma/\overline{\Sigma})$ is the logarithmic surface density, $\sigma_{\eta}$ is the standard deviation of the logarithmic surface density, and $\overline{\eta}=-1/2\sigma_{\eta}^{2}$ is its mean in analogy to $s$, $\sigma_{s}$, and $\overline{s}$ for the volume density PDF (Eq. 10). We use $\eta$ to denote the logarithmic surface density for notational consistency with other work and rely on context to disambiguate the logarithmic surface density and outflow efficiency.
@@
Using that relationship, Seon (2012) show that the relationship between the two dispersions is well-approximated by a constant of proportionality $\sigma_{\eta}^{2}=A\sigma_{s}^{2}$, where $A$ depends only on the forcing parameter. For $b_{\rm turb}=1/3,0.5,1$, they find best fit values $A=0.2,0.24,0.38$ respectively. We fit these data with a linear function using a least-squares regression, which combined with Equation 12, yields
@@
Using the column density PDF, we can compute the effective optical depth through a cell. We assume a constant opacity $\kappa$ such that $\tau=\kappa\Sigma$. Then the optical depth PDF is $p_{A}(\tau)=\tau^{-1}p_{A}(\eta)$. Consider a radiation field with intensity $I_{0}$ incident on the face of a cell with area $\Delta x^{2}$. Then the incident flux $F_{0}=I_{0}\Delta x^{2}$. The flux out the other side of the cell has contributions from each column
@@
where $\overline{\tau}=\kappa\overline{\Sigma}$ is the mean optical depth. Let $F=F_{0}e^{-\tau_{\rm eff}}$ define the effective optical depth. Then we have
[…]
In the limit $\mathcal{M}_{\rm turb}\ll 1$, the logarithmic surface density PDF becomes a delta function and the effective optical depth approaches the mean optical depth as expected.
The integral can be evaluated numerically using Gaussian quadrature as implemented in scipy.integrate.quad. For multiple simultaneous evaluations, we use the vectorized version implemented in scipy.integrate.quad_vec. In Figure 21, we show the effective optical depth relative to the mean optical depth as a function of mean optical depth for various turbulent Mach numbers and turbulence forcing parameters. At high Mach numbers and high mean optical depths, the effective optical depth is significantly reduced relative to the mean.
Figure 21: Effective optical depth (Eq. 75) in a turbulent medium relative to mean optical depth as a function of mean optical depth. Curves for different turbulent Mach numbers are shown in different colors indicated by the colorbar. Curves for compressive (resp. solenoidal) forcing are shown in solid (resp. dashed) lines. At high Mach numbers and high mean optical depths, the effective optical depth is significantly reduced relative to the mean.